### v1

# Representing 3D Shapes With 64 Latent Vectors for 3D Diffusion Models

Abstract
Constructing a compressed latent space through a variational autoencoder (VAE) is the key for efficient 3D diffusion models. This paper introduces COD-VAE, a VAE that encodes 3D shapes into a COmpact set of 1D latent vectors without sacrificing quality. COD-VAE introduces a two-stage autoencoder scheme to improve compression and decoding efficiency. First, our encoder block progressively compresses point clouds into compact latent vectors via intermediate point patches. Second, our triplane-based decoder reconstructs dense triplanes from latent vectors instead of directly decoding neural fields, significantly reducing computational overhead of neural fields decoding. Finally, we propose uncertainty-guided token pruning, which allocates resources adaptively by skipping computations in simpler regions and improves the decoder efficiency. Experimental results demonstrate that COD-VAE achieves 16$\times$ compression compared to the baseline while maintaining quality. This enables 20.8×20.8$\times$20.8 × speedup in generation, highlighting that a large number of latent vectors is not a prerequisite for high-quality reconstruction and generation.
1 Introduction
Learning a compressed latent space through a variational autoencoder [19] (VAE) has become a crucial component of recent diffusion models. Due to their iterative inference procedures, generating contents directly in the observation space (e.g., pixels) is computationally intractable. To address this challenge, most state-of-the-art generative models adopt a latent diffusion framework: a VAE first compresses the input data into a lower-dimensional latent space, and diffusion models are trained within this latent space.
Constructing a compact, well-structured latent space is even more crucial in the 3D domain. 3D objects are inherently irregular, sparse and continuous, which makes direct generation in the 3D domain more challenging. Previous research has explored various explicit 3D representations, e.g., points [43], voxels [37], and octrees [44]. These explicit representations typically demand larger latent sizes and specifically designed networks to process, making the training and the scaling of generative models difficult.
Figure 1: (top) Reconstruction IoU (left) and Rendering-FID of the generation results (right) with varying numbers of latent vectors ($M$). Our COD-VAE outperforms VecSet using 8$\times$ fewer latent vectors, achieving 20.8×20.8$\times$20.8 × generation speedup. (bottom) VecSet with $M=64$ struggles to capture details, while our model accurately reconstructs detailed and complex shapes of the objects.
Recently, 3DShape2VecSet [50] (VecSet) established a foundational framework for encoding 3D shapes into a set of 1D latent vectors. These 1D vectors provide a more compact latent space than the explicit representations, and are highly compatible with modern transformer implementations (e.g., FlashAttention [8]). Due to their simplicity and effectiveness, 1D latent vectors have been actively extended in subsequent research [52, 33, 53, 2, 10, 48, 6, 54].
The core component of VecSet is the cross-attention layer, which directly projects an input point cloud into 1D latent vectors and maps them back to continuous neural fields. These direct mappings, operating as learnable downsampling and interpolation, leads to the latent vectors only with a moderate compression ratio. Due to the limited compression ratio, VecSet still requires sufficiently many latent vectors for high-quality reconstruction, which results in excessive computational costs of 3D generative models. In addition, directly mapping latent vectors to neural fields through cross-attention further increases computational demands, creating another computational bottleneck.
In this paper, we introduce COD-VAE, a VAE that encodes 3D shapes into a COmpact set of 1D vectors with an improved compression ratio. COD-VAE replaces the direct mappings between points and latent vectors with a two-stage autoencoder scheme. This scheme enables our model to construct a significantly compressed latent space, thereby accelerating subsequent diffusion models. It achieves high-quality reconstruction with 16×16$\times$16 × fewer latent vectors, as well as efficiency improvement in nerual fields decoding.
The two-stage autoencoder scheme leverages intermediate representation spaces with a moderate compression ratio. These intermediate representations serve as bridges between 3D points and latent vectors. Concretely, our COD-VAE incorporates the following core components:
  • Our encoder introduces intermediate point patches, which are obtained by leveraging the attention-based downsampling of VecSet as a learnable point patchifier. Each encoder block progressively transforms high-resolution points into compact latent vectors by first mapping them into intermediate point patches and then further compressing them into latent vectors. The global information of the latent vectors are mapped back to points at the end of the block, further refining high-resolution features.
  • Our decoder leverages triplanes [3, 11] as intermediate representations for efficient yet effective decoding process. Inspired by [47], we treat dense triplane embeddings as mask tokens, and reconstruct them from the latent vectors using transformers. This design significantly reduces computational overheads of the neural fields decoding while also improving the reconstruction quality.
  • Apart from the efficient neural fields decoding, processing dense triplanes through transformers requires excessive computational resources. To mitigate this, we propose uncertainty-guided token pruning that reduces computations in simple regions, thereby achieving further efficiency improvement. Our decoder employs an auxiliary uncertainty head at the beginning, trained to predict reconstruction errors. By pruning regions with lower uncertainty, our decoder prioritizes computational resources for reconstructing more complex regions.
The compact latent space constructed by COD-VAE enables efficient and high-quality 3D generation. Experimental results on ShapeNet [4], and Objaverse [9] verify the effectiveness of our method, achieving 20.8×20.8$\times$20.8 × speedup in generation without sacrificing performance (see Figure 1). These results highlight that a large number of latent vectors is not a prerequisite for high-quality reconstruction and generation. We only need 64 latent vectors to surpass the results of VecSet with 512, or even 1024 latent vectors.
2 Related work
3D generative models.  Previous methods for 3D generation can be categorized into two main approaches. The first approach employs an optimization-based pipeline with 2D diffusion priors to generate 3D objects [34, 25, 23, 38, 42, 36]. These methods incur extensive computational costs due to the optimization process per object. Conversely, the second approach directly trains generative models on 3D datasets, which have shown promising results with significantly faster inference times. These native 3D generative models have widely exlpored, including GAN-based models [1, 40, 15, 3, 12, 32], autoregressive models [41, 30, 27, 45], and normalizing flows [46, 20, 18]. Recent breakthroughs of diffusion models have also inspired a wave of research in the 3D domain [31, 56, 39, 22, 16, 28], which typically adopt the latent diffusion framework [29].
3D VAEs.  Several studies have explored 3D VAEs with various latent representations. One line of works employs dense structures to model the latent space [7, 21, 24], which lead to quadratic [21, 24] or cubic [7] computational costs as the resolution increases. Other works have designed a more efficient latent space by leveraging sparse structures, such as points [43], octrees [44], sparse voxels [37], or irregular latent grids [49]. However, processing sparse structures is challenging and often requires specialized architectures and generation process. This makes the training of generative models difficult, limiting the scalability of these methods.
1D latent vectors.  VecSet [50] proposes a framework to model the latent space through 1D vectors. Its simplicity, compactness, and effectiveness has inspired a wide range of research [52, 10, 33, 53, 48, 6, 54]. The major challenge present in this line of works is the number of latent vectors required for high-quality reconstruction. As the direct mappings through the cross-attention lacks the capability to obtain higher compression ratio, successive generative models suffer from huge computation costs to generate these large sets of latent vectors. Additionally, decoding neural fields through the direct mapping causes another excessive computation costs. Our COD-VAE addresses these challenges with a two-stage autoencoder scheme, in which we utilize intermediate representation spaces as bridges.
Compact latent space.  The high computational costs of self-attention has led to efforts in the 2D domain for enhancing latent space compression [47, 5]. TiTok [47] proposes to represent images with highly compressed 1D latent tokens, which are obtained using transformers. DC-AE [5] achieves highly compressed 2D latent space to accelerate diffusion models. We aim to extend this paradigm to the 3D domain, by leveraging learnable point patches and triplanes as intermediate representations. Concurrent to our work, other studies [13, 51] introduce variational tokenizers with next scale prediction for 3D autogressive modeling.
Figure 2: Overview of COD-VAE. We propose a two-stage autoencoder scheme to obtain compact 1D latent vectors. The encoder leverages intermediate point patches with a moderate compression ratio, and the decoder reconstructs triplanes from the latent vectors.
3 Method
3.1 Preliminary of VecSet
VecSet [50] is a transformer-based VAE that directly maps points clouds into a set of 1D latent vectors and decodes them into continuous occupancy values. Given an input point cloud $\mathcal{P}=\{\mathbf{p}_{i}\}_{i=1}^{N}$ sampled from the object’s surface, VecSet encodes the point cloud into a set of 1D feature vectors =$\mathcal{F}=\{\mathbf{f}_{i}\in\mathbb{R}^{C}\}_{i=1}^{M}$ using the cross-attention:
$\mathcal{F}=\mathrm{CrossAttn}(\gamma(\hat{\mathcal{P}}),\ \gamma(\mathcal{P})),$ (1)
where $\gamma(\cdot)$ is a learnable positional embedding, and $\hat{\mathcal{P}}$ is the initial positions of $\mathcal{F}$, obtained by downsampling $\mathcal{P}$ with the farthest-point-sampling (FPS). Then the 1D latent vectors $\mathcal{Z}=\{\mathbf{z}_{i}\in\mathbb{R}^{D}\}_{i=1}^{M}$, where $D\ll C$, are obtained by compressing $\mathcal{F}$ along the channel dimension:
$\mathbf{z}_{i}=\mathrm{FC}_{\mu}(\mathbf{f}_{i})+\mathrm{FC}_{\sigma}(\mathbf{f}_{i})\cdot\epsilon,$ (2)
where $\mathrm{FC}_{\mu}(\cdot)$, $\mathrm{FC}_{\sigma}(\cdot)$ are linear layers that project the features into $D$-dimensional mean and variance, and $\epsilon\sim N(0,I)$. The resulting $\mathcal{Z}$ is a set of multivariate 1D Gaussian features, which is also regularized by KL divergence.
The decoder of VecSet first projects $\mathcal{Z}$ into $C$-dimensional features $\hat{\mathcal{F}}$, and processes them with several self-attention layers. Finally, a cross-attention layer maps these vectors into continuous occupancy values. For a query point $\textbf{q}\in\mathbb{R}^{3}$, the decoding process can be described as
$o(\mathbf{q})=\mathrm{FC}_{o}(\mathrm{CrossAttn}(\gamma(\mathbf{q})),\mathrm{SelfAttn}^{L}(\hat{\mathcal{F}})),$ (3)
where $\mathrm{SelfAttn}^{L}(\cdot)$ is a stack of $L$ self-attention layers, and $\mathrm{FC}_{o}:\mathbb{R}^{C}\rightarrow\mathbb{R}$ is a linear layer.
The cross-attention layers, which directly map points to latent vectors and vice versa, form the core of the VecSet framework. However, due to their limited compression capability, VecSet struggles to achieve latent vectors over a certain compression ratio. As a result, VecSet-based methods need a large number of latent vectors to obtain high-quality results, which yields substantial computational costs of diffusion models. Furthermore, directly decoding latent vectors into neural fields introduces an additional bottleneck, as neural fields decoding involves processing of enormous query points (i.e., over 2M points for $128^{3}$ grids).
3.2 COD-VAE architecture
We introduce COD-VAE, a model that represents 3D shapes using a set of compact 1D latent vectors. It replaces direct mappings of VecSet with a two-stage autoencoder scheme. This allows our model to construct the compressed latent space while enabling efficient neural fields decoding.
The overall architecture of COD-VAE is illustrated in Figure 2. Given an input point cloud $\mathcal{P}=\{\mathbf{p}_{i}\}_{i}^{N}$, our model encodes 1D latent vectors $\mathcal{Z}=\{\mathbf{z}_{i}\}_{i=1}^{M}$ from this point cloud. In the decoding stage, our latent decoder first decompress $\mathcal{Z}$ along the channel dimension, which is then decoded to dense triplane patches. Finally, we compute continuous occupancy values using the reconstructed triplanes and a shallow MLP. Our uncertainty-guided pruning improves the efficiency of the decoder. The following paragraphs describe details of each component.
Encoder.  Our encoder begins by obtaining initial positions of intermediate point patches and latent vectors. Same as VecSet, these are sampled from the input point cloud using FPS. The initial $C$-dimensional features of high-resolution points $\mathcal{G}^{(0)}=\{\mathbf{g}_{i}\}_{i=1}^{N}$, intermediate point patches $\mathcal{H}^{(0)}=\{\mathbf{h}_{i}\}_{i=1}^{L}$, and compact vectors =$\mathcal{F}=\{\mathbf{f}_{i}\ \}_{i=1}^{M}$, are computed by applying positional embedding $\gamma(\cdot)$ to their positions. Note that the sizes of the sets are $M\ll L\ll N$.
Figure 3 illustrates the design of our encoder block. Our encoder block progressive converts high-resolution point features to compact feature vectors through the intermediate point patches. It first aggregates high-resolution features $\mathcal{G}$ to $\mathcal{H}$ through the cross-attention, and process the aggregated patches via self-attention layers. The processed patches are then mapped to $\mathcal{F}$. Finally, the high-resolution features $\mathcal{G}^{(l)}$ are further enriched by projecting $\mathcal{F}$ onto them. The outputs of $l$-th encoder block, $\mathcal{H}^{(l)}$, $\mathcal{F}^{(l)}$, $\mathcal{G}^{(l)}$, can be expressed as
$\mathcal{H}^{(l)}=\mathrm{SelfAttn}^{3}(\mathrm{CrossAttn}(\mathcal{H}^{(l-1)},\mathcal{G}^{(l-1)})),$ (4)
$\mathcal{F}^{(l)}=\mathrm{SelfAttn}(\mathrm{CrossAttn}(\mathcal{F}^{(l-1)},\mathcal{H}^{(l)})),$ (5)
$\mathcal{G}^{(l)}=\mathrm{CrossAttn}(\mathcal{G}^{(l-1)},\mathcal{F}^{(l)}),$ (6)
where $\mathrm{SelfAttn}^{3}(\cdot)$ is a stack of 3 self-attention layers.
After the encoder blocks, the KL block transforms the compact feature vectors into 1D latent vectors. Similar to Eq. (2), it further compresses $\mathcal{F}$ along the channel dimension. This process yields a set of compact latent vectors, $\mathcal{Z}=\{\mathbf{z}_{i}\in\mathbb{R}^{D}\}_{i=1}^{M}$, where $D\ll C$. These latent vectors are also regularized by KL divergence, same as VecSet.
Figure 3: Encoder block. Our encoder block first projects high-resolution point features to the intermediate point patches and process them with self-attention layers. The processed patches are compressed into the compact 1D vectors.
Latent decoder.  Our decoding process begins with the latent decoder, which decompresses $\mathcal{Z}$ along the channel dimension. This process involves a linear layer that projects $\mathcal{Z}$ into $D$-dimensional features $\mathcal{F}$, and a stack of self-attention layers to process these features. Note that our latent decoder typically incurs negligible computation costs, as the size of our latent vectors set is small. The latent decoder is trained separately from the other components, with explicit objectives focusing on channel decompression. We provide further details of the two-stage training in Sec. 3.3.
Decoder.  After the latent decoding, our decoder reconstructs dense triplane features from the decompressed vectors $\mathcal{F}^{\prime}$ using transformer blocks. While these intermediate triplanes enable effective and efficient neural fields decoding, they also increase computational costs of the decoder, scaling quadratic to the triplane resolutions. To mitigate this, we introduce an uncertainty-guided token pruning that eliminates redundant computations in simple regions.
The architecture of our decoder is presented in Figure 2. We use a learnable token sequence $\mathbf{e}\in\mathbb{R}^{(\frac{R}{f}\times\frac{R}{f})\times C}$ to represent the positions of triplane tokens, where $R$ is the triplane resolutions and $f$ is the patch size. We initialize the triplane tokens by querying from $\mathcal{F}^{\prime}$ using a cross-attention block. Then the auxiliary head predicts uncertainty values of triplane tokens. We retain only the top $25\%$ tokens with the highest uncertainty values and prune the rest. The remaining tokens tokens are added with a shared mask token, concatenated with $\mathcal{F}^{\prime}$, and processed by ViT-style transformer blocks. Finally, we gather the processed triplane tokens and the pruned ones, which are projected by a linear layer to reconstruct full triplane features.
Neural fields decoding.  To reconstruct 3D objects, we decode continuous neural fields with the reconstructed triplane. For a query point $\mathbf{q}$, we retrieve features from each plane and add these features to obtain a query-wise feature. This feature is passed through a shallow MLP and converted to an occupancy value. Our triplane-based neural fields decoding process achieves substantial improvement in efficiency compared to VecSet, as it only requires computations for bilinear interpolation and a shallow MLP.
Figure 4: Decoder design. Our decoder prunes most of the triplane tokens at the beginning based on the predicted uncertainty values. The processed and pruned tokens are gathered at the last stage of the decoder to reconstruct the full triplane features.
3.3 VAE training
We train our COD-VAE using binary cross entropy loss, with KL divergence loss to regularize latent vectors and uncertainty loss to optimize the uncertainty head. Please refer to the supplements for the full objective of our model.
Uncertainty head.  The uncertainty head is trained to predict reconstruction error, which is computed using the initial triplane tokens. We first apply a sigmoid activation to the predicted uncertainty values and arrange them into triplanes. Then, we compute the binary cross entropy loss $\mathcal{L}_{rec}(\mathbf{q})$ for a query point $\mathbf{q}$. Following [11], we retrieve uncertainty values from the triplanes and compute the query-wise uncertainty $u(\mathbf{q})$ as
$u(\mathbf{q})=\psi_{xy}(U_{xy},\mathbf{q})\cdot\psi_{yz}(U_{yz},\mathbf{q})\cdot\psi_{xz}(U_{xz},\mathbf{q}),$ (7)
where $\psi_{C}$ projects $\mathbf{q}$ onto $C$-plane and retrieves the corresponding feature, and $U_{C}\in\mathbb{R}^{\frac{R}{f}\times\frac{R}{f}}$ is the triplane uncertainty on the $C$-plane. The uncertainty head is trained by minimizing the mean-squared-error (MSE) between $u(\mathbf{q})$ and $\mathcal{L}_{rec}(\mathbf{q})$. This objective encourages the uncertainty head to identify tokens that have higher reconstruction errors and need more computations. Our uncertainty-guided pruning eliminates redundant computations in simple regions, thereby addressing computation costs of the decoder.
Two-stage training.  Instead of jointly training all components, we adopt a two-stage training strategy. In the first stage, we train the autoencoder, a model without the KL block and the latent decoder. In the second stage, we freeze the trained components and optimize the KL block and the latent decoder by minimizing the MSE loss between the outputs of the latent decoder, $\mathcal{F}^{\prime}$, and the encoder outputs, $\mathcal{F}$. We also minimize the KL divergence loss, as well as the reconstruction error for stable training in the second stage. This two-stage training guides the latent decoder to focus on the channel decompression of the latent vectors.
3.4 Diffusion model
We train a latent diffusion model on the compact latent space constructed by our COD-VAE. Following VecSet [50], we adopt transformer-based denoising networks and use the diffusion formulations from EDM [17]. While we do not explicitly aim to improve the generation process itself, the compact latent space provided by COD-VAE significantly enhances the efficiency of our diffusion models.
4 Experiments
To verify the effectiveness of our method, we first evaluate the reconstruction quality of our COD-VAE. Then, we assess the results of our diffusion models on both class-conditioned and unconditional generation. We provide more details on the experimental setups in the supplements.
Datasets.  We conduct the experiments on two public datasets, ShapeNet-v2 [16] and Objaverse [9]. ShapeNet [4] consists of 3D CAD models from 55 categories. We follow the experimental setups of [49, 50], and use same train/validation splits and preprocessed meshes.
Objaverse contains around 80K objects, which are more diverse and complex than the objects in ShapeNet. Due to its large size, we sample the subset of around 150K objects based on the list provided in [36]. We split the dataset into 135,511 objects for training and 14,663 for validation. Following [48], we employ ManifoldPlus [14] to make watertight meshes, and normalize them to be in unit cubes. We use the full dataset for training VAE, and use the subset with around 90K objects for training diffusion models.
Baselines.  We select VecSet [50] as our primary baseline and report its results for various numbers of latent vectors $M$. We also compare with additional state-of-the-art methods, including 3DILG [49], GEM3D [33]. We reproduce results of VecSet using its official code, with the latent channel dimension $D=32$. We report the results of other baselines using either the provided pretrained weights, or their official code when the weights are unavailable.
Evaluation metrics.  For the reconstruction experiments, we follow VecSet and use the mean intersection-over-union (IoU) of 50K query points uniformly sampled from unit cubes. We also report Chamfer Distance (CD) and F-score (F1), computed from 10K points sampled from both the reconstructed and ground truth meshes. For the generation experiments, we adopt FID-based scores as our main metrics due to their reliability. Following [55, 50, 33], we measure the FID of the images rendered from 20 viewpoints (Rendering-FID), as well as the Fréchet PointNet++ Distance using features from PointNet++ (Surface-FPD). To compute FPD, we train PointNet++ [35] on ShapeNet-55 classification and use this model to extract features. We additionally measure MMD, COV, and 1-NNA, following [55]. MMD assess the fidelity, COV measures the coverage, and 1-NNA evaluates the diversity of the generated shapes. We measure the throughput using a single A6000 GPU.
Implementation details.  Our model consists of 4 encoder blocks and 12 decoder layers, with $C=512$ and $D=32$. During training, we sample 4,096 query points uniformly from unit cubes and 4,096 points near the objects’ surfaces. We train the autoencoders for 1,600 epochs on ShapeNet and 300 epochs on Objaverse. The second-stage VAE training follows the same number of epochs. To train our models, we use 4 RTX 4090 GPUs for ShapeNet and 4 A6000 GPUs for Objaverse. Training VAEs takes about 3 days for ShapeNet and 6 days for Objaverse, and training diffusion models takes less than 4 days for both ShapeNet and Objaverse. Our code will be made publicly available.
Figure 5: Qualitative reconstruction results. We deliver the reconstruction results of VAEs.
4.1 Reconstruction results
For ShapeNet experiments, we follow previous research [49, 50] and report results of both autoencoders (without channel compression by KL block) and VAEs. We also set the input point cloud size $N=2048$, the test grid resolution to 128. For our models, we set $L=512$, and the triplane resolution to 128. For Objaverse, due to the complexity of the objects, we increase the input point cloud size to $N=8192$. We increase the number of latent vectors of VecSet by 2 as well as the number of our point patches ($M=L=1024$), and the triplane resolution to 256.
Autoencoder VAE
[…]
3DILG (M=512) 95.9 0.013 98.0 94.7 0.013 97.6
VecSet (M=32) 87.8 0.021 91.3 87.0 0.021 90.6
VecSet (M=64) 91.2 0.017 94.7 90.9 0.017 94.4
VecSet (M=512) 96.3 0.013 98.0 96.2 0.013 98.0
Ours (M=32) 96.1 0.012 98.0 95.8 0.013 97.8
Ours (M=64) 96.5 0.012 98.2 96.3 0.012 98.0
Table 1: Reconstruction results on ShapeNet.
Results on ShapeNet.  Table 1 reports the reconstruction results on ShapeNet. Our models, both autoencoders and VAEs, outperform all other methods with only using 64 latent vectors. Even with 32 latent vectors, our models achieve reconstruction quality comparble to VecSet ($M=512$), offering a more efficient yet effective option for diffusion models. In contrast, simply reducing the number of latent vectors of VecSet leads to noticeable degradation in both autoencoders and VAEs. This is also evident in the qualitative results presented in Figure 5 (a). Thanks to the two-stage autoencoding, our models with $M=64$ achieve high-quality reconstruction results, whereas VecSet with $M=64$ fails to represent fine details of the objects.
Results on challenging objects.  We then evaluate the performance on Objaverse, which comprises more challenging and diverse objects than ShapeNet. As reported in Table 2, our model with $M=64$ exhibits even higher reconstruction quality than VecSet with $M=1024$. Our VAE also significantly boosts the efficiency of the generation, achieving 21.6×21.6$\times$21.6 × higher throughput than VecSet with $M=1024$. We also deliver the qualitative results in Figure 5 (b). Reducing the number of latent vectors of VecSet leads to the severe degradation. These results showcase that we can represent 3D shapes in high-quality using 64 latent vectors.
4.2 Generation results
To further validate our latent representations, we evaluate the generation performance of diffusion models trained on our latent space. We deliver the results of class-conditioned and unconditional generation, which are conducted on ShapeNet and Objaverse, respectively. We follow VecSet and set the number of sampling steps 18 for ShapeNet. For Objaverse, we increase the sampling steps to 50 due to the complexity of this dataset.
VAE Generation
[…]
VecSet ($M=64$) 71.2 0.054 85.2 1.99 5.46
VecSet ($M=1024$) 79.8 0.051 87.0 0.23 9.96
Ours ($M=64$) 79.9 0.046 88.0 4.97 3.08
Table 2: Reconstruction results on Objaverse. The GPU memory (GB) is measured with a batch size 16.
Method Rendering-FID $\downarrow$ Surface-FPD $\downarrow$ MMD $\downarrow$ COV (%) $\uparrow$ 1-NNA (%) $\downarrow$ Throughput (sample/s) $\uparrow$
CD EMD CD EMD CD EMD Sampling Full
3DILG [49] 64.83 2.585 6.270 10.98 58.16 62.80 61.49 59.90 0.14 0.09
GEM3D [33] 47.98 0.997 5.381 10.41 66.64 66.05 54.74 56.02 0.09 0.07
VecSet ($M=32$) [50] 65.56 0.800 4.999 10.44 84.51 82.85 54.32 54.76 39.77 2.81
VecSet ($M=64$) 54.47 0.629 5.090 10.49 84.89 84.45 55.04 54.94 21.92 2.63
VecSet ($M=512$) 44.18 0.521 4.807 10.21 85.20 84.23 53.78 55.06 2.59 1.16
Ours ($M=32$) 37.34 0.473 5.020 10.38 84.96 84.23 53.11 54.09 41.19 24.17
Ours ($M=64$) 37.05 0.460 5.025 10.38 86.03 85.37 54.89 54.07 22.49 16.09
Table 3: Class-conditioned generation results on ShapeNet. The scales of MMD are $10^{-3}$, and $10^{-2}$ for CD, and EMD, respectively. The ‘Sampling’ denotes the throughput of the latent sampling, and the ‘Full’ denotes the throughput of the entire process.
Batch size 32 Batch size 64
Speed Memory Speed Memory
VecSet ($M=512$) 1.01 39.01 - OOM
Ours ($M=32$) 2.60 5.17 1.37 8.33
Ours ($M=64$) 2.23 5.63 1.20 9.23
Table 4: Training efficiency of the diffusion models. We report the speed (iter/s) and the GPU memory (GB) on ShapeNet.
Figure 6: Qualitative class-conditioned generation results on ShapeNet.
Category-conditioned generation.  Table 3 reports the class-conditioned generation results. The results are averaged over 5 commonly used classes, which are airplane, car, chair, table, and rifle. Our models achieve the highest performance in FID-based scores (Rendering-FID, Surface-FPD) and outperform other methods in most additional metrics. In contrast, VecSet with $M=32,64$ delivers degraded results, as reflected in FID-based scores and further shown in Figure 6. Notably, our method significantly improves the efficiency of generative models, achieving 13.9×13.9$\times$13.9 × and 20.84×20.84$\times$20.84 × improved throughput. The compact latent space also improves the training efficiency, as reported in Table 4.
Figure 7: Unconditional generation results of our model. We present the results with $M=64$.
Unconditional generation.  We present that our diffusion model, trained on the compact latent space, can learn the distributions of a wide range of objects. We deliver the qualitative unconditional generation results of our model in Figure 7. Our model generates diverse types of shapes in high-quality with significantly improved efficiency.
4.3 Ablation study
To verify the effectiveness of the proposed components, we conduct ablation studies on ShapeNet reconstruction. We report the results of autoencoders with $M=32$.
Encoder design.  We first replace our encoder with the encoder of VecSet, which utilizes the cross-attention to directly map point features to latent vectors and self-attention layers to process latent vectors. We compare encoders with various number of cross-attention layers. As shown in Table 5 (a), our encoder outperforms the models with VecSet’s encoders by a large margin. Using more cross-attention layers does not improve performance beyond a certain point. This highlights the importance of intermediate point patches, which enables higher compression ratio by proressive transformations.
Decoder design.  We also replace our decoder with the cross-attention layer as in VecSet, which directly maps latent vectors to query points. We report the ablation results on the decoder design in Table 5 (b). Our decoder improves the overall reconstruction quality as well as efficiency, compared to that of VecSet. In addition, the intermediate triplane representations significantly reduces computation costs of the neural fields decoding, which accompanies more than 20M query points to reconstruct $128^{3}$ grids. Adding more cross-attention layers could offer better reconstruction quality, but at the cost of excessive computations.
[…]
CrossAttn 91.2 0.016 94.7
CrossAttn ×$\times 12$ 93.4 0.014 96.5
CrossAttn ×$\times 24$ 93.5 0.014 96.6
Ours 96.1 0.012 98.0
(a) Encoder design.
[…]
(sample/s)$\downarrow$
CrossAttn 95.8 0.013 97.8 3.43
CrossAttn×$\times 3$ 96.3 0.013 98.0 1.47
Triplane 96.1 0.012 98.0 42.68
(b) Decoder design.
[…]
Prune 0% 96.2 0.012 98.0
Prune 50% 96.2 0.012 98.0
Prune 75% 96.1 0.012 98.0
Prune 90% 95.9 0.012 98.0
(c) Pruning ratio results with $R=128$.
Table 5: Quantitative ablation results with $M=32$ on ShapeNet reconstruction.
Pruning Speed (iter/s) Memory (GB)
ratio $R=128$ $R=256$ $R=128$ $R=256$
0% 4.21 1.17 6.79 26.99
50% 5.14 2.21 5.50 11.87
75% 5.60 3.09 5.02 7.80
90% 5.78 3.71 4.78 6.33
Table 6: Training efficiency analysis of the pruning.
Pruning ratio.  Finally, we validate our uncertainty-guided pruning by changing the pruning ratio. As shown in Table 5 (c) and Table 6, training costs are effectively reduced as the pruning ratio increases with slight performance drops. With $R=128$, pruning 50% triplane tokens maintains the reconstruction quality while achieving 1.22×1.$2$ speedup and 0.81×0.81$\times$0.81 × memory consumption for training the autoencoders. Further increasing the pruning ratio to 75% results in a 1.33×1.$3$ speedup and 0.73×0.73$\times$0.73 × memory, with only a slight drop in performance, which we choose as our main setting. The benefits of our uncertainty-guided pruning become more evident as the triplane resolution increases to $R=256$. This improved efficiency allows our model to scale more effectively to datasets with more complex samples, e.g., Objaverse.
5 Discussion
Accelerating the query decoding.  Reconstructing neural fields typically requires evaluating a vast number of query point (over 2M points for $128^{3}$ test grids). While our triplane-based query decoding significantly reduces these overheads, it often hinders maximizing throughput (see Table 3, Ours with M=32). This can be further addressed by utilizing the improved query points sampling, e.g., multi-resolution approach as proposed in [26]. We further provide runtime analysis with this technique in the supplements.
Uncertainty visualization.  To predicted uncertainty values serve as the importance score of the triplane tokens. As presented in Figure 8, our uncertainty head assigns higher scores to more detailed regions, placing more computations on them. This allows our model to avoid heavy computations for simple parts, and place more computational resources on reconstructing details and thin structures.
Figure 8: Visualizations of the predicted uncertainty values.
Num. latents (M) 64 128 256 512 1024 2048
Memory (GB) 1.13 1.64 2.86 5.65 13.63 38.66
Max batch size 128 64 32 16 4 2
Throughput (sample/s) 652.45 319.00 149.74 65.16 24.95 8.37
Table 7: Running statistics of the denoising network with various $M$. The network has 24 layers with 512 channel. The GPU memory is measured with a batch size 2, and the throughput is measured with the corresponding maximum batch size.
Scaling and the compact latent space.  The efficiency improvement of the compact latent space becomes more apparent as the number of latent vectors increases, as shown in Table 7. With the same compression ratio 8×8$\times$8 ×, reducing $M$ from 512 to 64 leads to the 10.0×10.0$\times$10.0 × throughput and 0.20×0.20$\times$0.20 × memory, while reducing from 2048 to 256 results in the 17.9×17.9$\times$17.9 × throughput and 0.07×0.07$\times$0.07 × memory usage. We highlight that constructing the compact latent space can be a key for scaling generative models to more complex scenes, which usually requires a large number of latent vectors.
Limitations and future works.  While our COD-VAE successfully alleviates bottlenecks of the generation process, parameter counts of VAE and computations of the encoder are increased compared to VecSet (see the supplements for more details). We anticipate that enhancing the encoder efficiency can further boost up the training of VAEs and generative models. Moreover, our VAE has the potentials to encode information from multiple sources, e.g., images and videos. This can be an interesting avenue for future works.
6 Conclusion
Reducing the latent size is the key for efficient 3D generation. Our COD-VAE with a two-stage autoencoding scheme represents 3D shapes with 64 latent vectors without sacrificing reconstruction and generation quality. This compact latent vectors set enables efficient and high-quality generation, suggesting a new direction for 3D generative models.